# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v1

@@ II.2.3 Relating $\Psi^{a}$ to the dimension and timescales of activity
$\left\langle J_{ij}\right\rangle_{\bm{J}}=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle_{\bm{J}}=\frac{g^{2}}{N}.$ (14)
The magnitude of a typical coupling is $g/\sqrt{N}$. We refer to $g$ as the coupling strength. For the canonical dynamics of Eq. 2, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\rightarrow\infty$. In this paper we assume that the network is non-quiescent.
We compute $C^{a}(t_{1},t_{2})$ and $\Psi^{a}(\bm{\tau})$ in the limit $N\rightarrow\infty$ under the disorder average. We use the subscript $\star$ to denote such limiting, disorder-averaged values:
$C^{a}_{\star}(t_{1},t_{2})$ $=\lim_{N\rightarrow\infty}\left\langle C^{a}(t_{1},t_{2})\right\rangle_{\bm{J}},$ (15)
$\Psi^{a}_{\star}(\bm{\tau})$ $=\lim_{N\rightarrow\infty}\left\langle\Psi^{a}(\bm{\tau})\right\rangle_{\bm{J}}.$ (16)
These values also correspond to the saddle point of a path integral. Both of these functions are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network. In this paper, we are interested in statistically stationary, temporally fluctuating states such that $C^{\phi}_{\star}(t_{1},t_{2})=C^{\phi}_{\star}(t_{2}-t_{1})$.
@@
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within this single-site process, i.e., with respect to the Gaussian distribution of $\eta^{x}(t)$. Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily. This single-site problem can be derived through either a single-site cavity calculation (see [39]) or a saddle-point condition in a path integral (see [40, 41]). For the i.i.d. couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq. 18.
@@
[…]
@@
II.4 Definition of the random-mode model
Figure 3: Schematic of the random-mode model. Upper: Couplings $\bm{J}$ are generated as a sum of outer products, ${\bm{L}}_{a}{\bm{R}}^{T}_{a}$, with component strengths $D_{a}$. Lower: The two-point function $C^{\phi}_{\star}(\tau)$ and four-point function $\Psi^{\phi}_{\star}(\tau)$ are calculated in terms of the statistics of $D_{a}$. The two-point function depends only on the effective gain $g_{\text{eff}}$, while the four-point function depends on both $g_{\text{eff}}$ and $\text{PR}^{D}$, the effective dimension of the connectivity determined by the $D_{a}$ distribution.
Many studies have examined the dynamics of networks with i.i.d. couplings, as described in the previous section. However, the coupling matrices of real-world networks often have approximate low-rank spectral structure, deviating dramatically from the i.i.d. assumption. To analyze networks with such properties, we introduce the random-mode model. This model generates a coupling matrix $\bm{J}$ of the form
$\bm{J}=\sum_{a=1}^{M}D_{a}\bm{L}_{a}\bm{R}_{a}^{T},$ (22)
where $\bm{R}_{a}$ and $\bm{L}_{a}$ are “input” and “output” modes in the neuron space and $D_{a}>0$ is a weighting factor called the component strength. The effect of multiplying by $\bm{J}$ is to project a neuronal activity vector onto $\bm{R}_{a}$, scale it by $D_{a}$, project the result along $\bm{L}_{a}$, and sum these contributions for all $a=1,\ldots,M$ (Fig. 3, top). Eq. 22 can be written
$\bm{J}=\bm{L}\bm{D}\bm{R}^{T},$ (23)
where $\bm{L}$ and $\bm{R}$ are $N\times M$ matrices and $\bm{D}=\text{diag}(D_{1},D_{2},\ldots,D_{M})$.
A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$. We consider the limit where both $M$ and $N$ approach infinity while maintaining a fixed ratio,
$\alpha=\frac{M}{N}=\mathcal{O}(1).$ (24)
This scaling results in extensive-rank connectivity, meaning that the rank of $\bm{J}$ grows linearly with $N$. Importantly, choosing $\alpha$ to be small (or choosing the component strengths to decay rapidly; see below) allows for connectivity that is low-rank compared to the number of neurons.
For now, we assume that the matrices $\bm{L}$ and $\bm{R}$ have independent and identically distributed elements with zero mean and variance $1/N$. We assume these elements are Gaussian, although our results should not depend on the exact form of the single-element density for large $N$ due to the central limit theorem.
A key feature of the random-mode model is that we have full control over the component strengths $D_{a}$, which are analogous to the singular values of $\bm{J}$ (we make this link precise in Sec. II.4.1). We assume only that these component strengths follow a distribution that remains constant as $N$ grows, ensuring a well-defined large-$N$ limit. By varying this distribution, we can explore the impact of the spectral structure of the connectivity on collective features of network activity (Fig. 3, bottom).
We denote the $n$-th moment of the $D_{a}$ distribution by
$r_{n}=\left\langle D^{n}\right\rangle_{D}.$ (25)
It will be useful to define the participation ratio,
$\text{PR}^{D}=\frac{r_{2}^{2}}{r_{4}},$ (26)
from which we define the “effective rank” of the couplings,
effective $\text{effective rank}=\alpha\text{PR}^{D}.$ (27)
This quantity is related to the participation ratio of the singular value spectrum of $\bm{J}$ via Eq. 28 below. We will show that this effective rank is a key feature of $\bm{J}$ that controls the collective activity structure.
II.4.1 Relationship to singular value decomposition
The random-mode model has similarities to the SVD, which decomposes a rank-$M$ matrix $\bm{J}$ as $\bm{U}\bm{S}\bm{V}^{T}$. Here, $\bm{U}$ and $\bm{V}$ are $N\times M$ matrices whose columns are the left and right singular vectors (or modes), respectively, and $\bm{S}=\text{diag}(S_{1},S_{2},\ldots,S_{M})$ is a diagonal matrix containing the singular values, $S_{a}>0$. The left singular vectors are orthonormal among themselves, as are the right singular vectors (although the left and right singular vectors are generally not orthonormal to each other). In contrast, the columns of $\bm{L}$ and $\bm{R}$ in the random-mode model have i.i.d. statistics, resulting in random, $\mathcal{O}(1/\sqrt{N})$ overlaps among both within- and across-matrix columns. This i.i.d. assumption offers analytical tractability by eliminating the need to enforce orthonormality, a complicated global constraint (but see [43], who handle random orthonormal matrices in a regression setting). Furthermore, by not explicitly modeling overlaps between $\bm{L}$ and $\bm{R}$, the random-mode model parameterizes $\bm{J}$ using only the $\mathcal{O}(N)$ component strengths $D_{a}$, although such overlaps can be introduced (see Sec. VII).
Given a matrix $\bm{J}$ drawn from the random-mode model, we can examine the relationship between the singular values, $S_{a}$, and the component strengths, ${D}_{a}$. The non-orthonormality of $\bm{L}$ and $\bm{R}$ introduces some discrepancy between the two sets of values. We first compare the participation ratio of the squared singular value spectrum, $\text{PR}^{S}$, to the effective rank, $\alpha\text{PR}^{D}$. We find:
$\text{PR}^{S}=\frac{\alpha\text{PR}^{D}}{1+2\alpha\text{PR}^{D}},$ (28)
where $\text{PR}^{S}=\langle S^{2}\rangle_{S}^{2}/\langle S^{4}\rangle_{S}$, and $S$ are the singular values of $\bm{J}$. This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$. For $\alpha\text{PR}^{D}\ll 1$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2}),$ demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime. In this regime, the random modes approximate an orthonormal set of vectors as $N\rightarrow\infty$ since the $\mathcal{O}(1/\sqrt{N})$ overlaps between modes can be neglected when there are few of them. While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, Eq. 28 can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
We further quantify the discrepancy between singular values and component strengths in the case where $D_{a}=1$ for all $a$. Using methods from free probability theory [44, 45], we compute that the $M$ singular values are distributed over a range with boundaries
$S_{\pm}=\sqrt{1+\frac{5\alpha}{2}-\frac{\alpha^{2}}{20}\pm\left(1+\frac{\alpha}{8}\right)^{\frac{3}{2}}\sqrt{8\alpha}}.$ (29)
This is in agreement with large-$N$ simulations (Fig. 4). For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha).$ While the random-mode model spreads the nonzero singular values over a range, this spreading becomes negligible for small $\alpha$. More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
II.4.2 Comparison to low-rank recurrent neural networks
A related parameterization of connectivity has been explored in low-rank recurrent neural networks, where the coupling matrix is expressed as a sum of outer products of input modes $\bm{n}_{a}$ and output modes $\bm{m}_{a}$:
$\bm{J}=\frac{1}{N}\sum_{a=1}^{R}\bm{m}_{a}\bm{n}_{a}^{T},$ (30)
with $R=\mathcal{O}(1)$ being the connectivity rank [30, 46, 31, 47]. In some instances, this low-rank component is added to an i.i.d. matrix modeling unstructured “background” connectivity [30, 46].
Several features separate this approach from the random-mode model. Low-rank recurrent neural networks maintain a constant rank $R$ as $N\rightarrow\infty$, yielding intensive-rank connectivity that models low-dimensional activity, whereas the random-mode model uses extensive-rank connectivity, modeling high-dimensional activity. Furthermore, low-rank recurrent neural networks allow for the most general possible parameterization of overlaps between modes, captured by a $R\times R$ overlap matrix $\bm{m}_{a}^{T}\bm{n}_{b}$. Specifying such overlaps in the random-mode model would require $\mathcal{O}(N^{2})$ parameters, comparable to defining the full coupling matrix. To reduce parameter complexity, the random-mode model specifies only the variances of mode components via the component strengths $D_{a}$, disregarding overlaps. This substantially reduces parameter count while preserving interesting network dynamics. In Sec. VII, we introduce $\mathcal{O}(N)$ additional parameters describing a restricted overlap structure among modes, specifying only diagonal elements of the overlap matrix. Further possible overlap forms are discussed in Sec. VIII.
The mean-field descriptions of network dynamics produced by these connectivities also differ markedly. In low-rank recurrent neural networks, low-dimensional network activity is described by a finite set of order parameters representing projections of network activity onto input modes $\bm{n}_{a}$, obeying effective nonlinear dynamics. Conversely, the extensive-rank connectivity in the random-mode model generates high-dimensional chaotic activity, described by covariance functions capturing temporal fluctuations. While the random-mode model’s mean-field analysis introduces latent variables given by projections of network activity onto input modes, their extensive number precludes their use as order parameters.
A final distinction lies in the scaling of the coupling matrix $\bm{J}$ and its implications for the alignment of activity with connectivity. In low-rank recurrent neural networks, elements of $\bm{J}$ scale as $\mathcal{O}(1/N)$ and the input-output mode overlaps are $\mathcal{O}(1)$, resulting in strong $\mathcal{O}(1)$ alignment of network activity with input modes. This is reminiscent of the rich regime of neural networks [48, 49, 50, 51]. Conversely, in the random-mode model, the elements of $\bm{J}$ scale as $\mathcal{O}(1/\sqrt{N})$ and the input-output mode overlaps are $\mathcal{O}(1/\sqrt{N})$, leading to weak $\mathcal{O}(1/\sqrt{N})$ alignment of network activity with any given input mode; the pre-activations are $\mathcal{O}(1)$ since each neuron combines input from $\mathcal{O}(N)$ modes. This is reminiscent of the lazy regime of neural networks [49, 52, 53].
@@ II.4.2 Comparison to low-rank recurrent neural networks
III Path-integral calculation of the two- and four-point functions
The path-integral approach provides a systematic framework for analyzing neural-network dynamics, allowing for controlled approximations such as $1/N$ expansions [54, 55, 41, 56]. In this section, we demonstrate how to formulate the calculation of the four-point function $\Psi^{a}_{\star}(\bm{\tau})$ within this formalism, allowing us to solve complex models more efficiently than the two-site cavity method by following a straightforward procedure whose steps we outline at the end of Sec. III.3.
III.1 Relating $\Psi^{a}$ to fluctuations of the time-by-time covariance
Figure 5: Neuron-time duality and its relation to the dimension of activity. Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d. couplings. The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$, given by Eq. 9. If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$ since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10]. However, based on the neuron-by-neuron definition, Eq. 7, $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity. This apparent contradiction is resolved by considering fluctuations around the saddle point. The necessity of these fluctuations becomes clear when viewed through the lens of neuron-time duality. The presence of nonzero pairwise covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system. Equivalently, this structure manifests as the system revisiting similar states in phase space over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig. 5). To capture this behavior, we define fluctuations around the saddle point as
$\delta C^{a}(t_{1},t_{2})=C^{a}(t_{1},t_{2})-C^{a}_{\star}(t_{2}-t_{1}).$ (31)
We express $\Psi^{a}(\bm{\tau})$ purely in terms of these fluctuations as
noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences. Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured by
$N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\star}\\ =\lim_{N\rightarrow\infty}N\left\langle\delta C^{a}(t_{1},t_{2})\delta C^{a}(t_{3},t_{4})\right\rangle_{\bm{J}}.$ (33)
Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we use the temporal limits
[…]
$t_{4}$ $=t_{2}+\tau_{2},$ $\tau_{1},\tau_{2}$ $\sim\mathcal{O}(1).$
III.2 Path-integral formalism and two-point functions
Given any fixed $\bm{J}$, we express the network dynamics using a path integral in the Martin–Siggia–Rose/Janssen–De Dominicis–Peliti formalism [57]:
$Z[\bm{J}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\left(i\sum_{i}\hat{x}_{i}T[{x}_{i}]-i\sum_{i,j}J_{ij}\hat{x}_{j}\phi_{j}\right),$ (35)
[…]
@@ III.2 Path-integral formalism and two-point functions
For simplicity, we begin with the i.i.d. model of $\bm{J}$ with mean zero and variance $g^{2}/N$. After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq. 5, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$. This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@
$W[C^{\phi},\hat{C}^{\phi}]=\int\mathcal{D}x\int\mathcal{D}\hat{x}\\ \times\exp\left(ixT[\hat{x}]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi\right).$ (38)
In the limit $N\rightarrow\infty$, the saddle point dominates this integral. The derivatives of the action, which are zero at the saddle point, are
@@
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$. This recovers the same single-site process (Eq. 17, Eq. 18) and self-consistency condition (Eq. 19) described in Sec. II.3.
III.3 Four-point functions
@@
$Z=\exp\left(-N\mathcal{S}[C^{\phi}_{\star},0]\right)\int\mathcal{D}\delta C^{\phi}\mathcal{D}\delta\hat{C}^{\phi}$ (44)
×$\times\exp\left(-\frac{N}{2}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}^{T}\begin{pmatrix}H_{C^{\phi}C^{\phi}}&H_{C^{\phi}\hat{C}^{\phi}}\\ H_{\hat{C}^{\phi}{C}^{\phi}}^{T}&H_{\hat{C}^{\phi}\hat{C}^{\phi}}\end{pmatrix}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}\right),$
@@
$N\left\langle\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})\right\rangle_{\star}$ (45)
$=\left[H_{C^{\phi}C^{\phi}}-H_{C^{\phi}\hat{C}^{\phi}}H_{\hat{C}^{\phi}\hat{C}^{\phi}}^{-1}H_{\hat{C}^{\phi}C^{\phi}}\right]^{-1}(t_{1},t_{2},t_{3},t_{4}).$
In principle, we have all the necessary components to evaluate this expression: the single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process. We could then take the temporal separation limits of Eq. 34 to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$. The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse. To circumvent this problem, we show in Appendix C that the temporal limits of Eq. 34 can be taken before taking the inverses. Under these temporal limits, the Hessian blocks are
@@
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (47)
$\hskip 40.00006pt=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(-\tau_{1})S^{\phi}_{\star}(-\tau_{2}),$
@@
where we used the fact that multiplication by $-i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$. Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space. In this representation, each Hessian block becomes a frequency-dependent scalar,
@@
$H_{C^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-1+g^{2}\left(S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})\right)^{*},$ (50)
@@
To summarize, we compute $\Psi_{\star}^{\phi}(\bm{\tau})$ through the following steps (in the next section, we show how to compute $\Psi^{x}_{\star}(\bm{\tau})$):
  1. 1.
    Write down the path integral, integrate out disorder, and introduce two-point fields and conjugate partners to factorize the action along extensive indices.
  2. 2.
    Sovle the saddle-point equations of the resulting action, yielding the covariance and response functions.
  3. 3.
    Derive the frequency-dependent Hessian by computing second derivatives of the action with respect to the two-point fields, applying the temporal limits of Eq. 34, and transforming to Fourier space. Then, determine the element of the inverse frequency-dependent Hessian corresponding to $\Psi^{\phi}_{\star}(\bm{\omega})$.
This approach allows for the efficient calculation of two- and four-point functions for a given model, requiring only a few lines of computation directly from the action of the theory. To demonstrate the flexibility of the path-integral approach in computing $\Psi^{a}_{\star}(\bm{\tau})$, we apply it to compute $\Psi^{\phi}_{\star}(\bm{\tau})$ for a generalization of the i.i.d. model in which there are correlations between reciprocal couplings, recovering the result in [25] from the two-site cavity method (Appendix E).
III.4 Adding sources

### v2

@@ Appendix D Alternative measures of effective dimensionality
$\left\langle J_{ij}\right\rangle=0,\hskip 15.00002pt\left\langle J_{ij}^{2}\right\rangle=\frac{g^{2}}{N}.$ (68)
The magnitude of a typical coupling is $g/\sqrt{N}$. We refer to $g$ as the coupling strength. For the canonical dynamics $T[x](t)=(1+\partial_{t})x(t)$, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\to\infty$. In this paper we assume that the network is non-quiescent.
@@
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within this single-site process, i.e., with respect to the Gaussian distribution of $\eta^{x}(t)$. Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily. This single-site problem can be derived through either a single-site cavity calculation (see [90]) or a saddle-point condition in a path integral (see [54, 55]). For the i.i.d. couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j=1}^{N}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq. 70.
@@
$+U(\bm{\omega})C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})+\text{h.c.},$ (73a)
@@
Appendix F Path-integral calculation of two- and four-point functions for i.i.d. couplings
Figure 9: Relationship between principal component timescales and two- and four-point correlation functions. (a) Example activity traces of principal components $p^{\phi}_{k}(t)$. (b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues. (c) Same as (b) but weighted by the squared eigenvalue. (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$. Dots: simulation results. Lines: theoretical predictions. All panels describe the same network of $N=4000$ neurons with i.i.d. couplings with variance $g^{2}/N$ with coupling strength $g=6$.
The general program, for i.i.d. or structured couplings, is as follows:
  1. 1.
    Formulate the field theory. Begin with the path integral $Z[\bm{J}]$ and average over the connectivity disorder. Introduce auxiliary fields representing two-point functions (such as $C^{\phi}(t_{1},t_{2})$) along with their conjugate partners. Use integral representations of delta functions to enforce field definitions, factorizing the exponential over extensive dimensions. This yields a statistical field theory with action $N$ times an order-one function of the auxiliary fields.
  2. 2.
    Solve the saddle-point equations. Find the saddle point by setting action derivatives to zero. This produces self-consistent equations for the two-point functions, completing the standard dynamical mean-field theory (DMFT) analysis.
  3. 3.
    Compute the Hessian. Calculate the time2-by-time2 Hessian matrix, which depends on four time variables $(t_{1},t_{2},t_{3},t_{4})$ and characterizes fluctuations around the saddle point.
  4. 4.
    Apply temporal separation limits. Impose the temporal limits (Eq. 34), making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$. Transform to Fourier space to obtain a low-dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
  5. 5.
    Invert the frequency-space Hessian. Perform matrix inversion to directly obtain $\Psi^{a}_{\star}(\bm{\omega})$. If needed, apply inverse Fourier transform to recover $\Psi^{a}_{\star}(\bm{\tau})$.
@@
Here we consider i.i.d. $\bm{J}$ with mean zero and variance $g^{2}/N$. After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq. 17, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$. This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@
$W[C^{\phi},\hat{C}^{\phi}]=\int\mathcal{D}x\int\mathcal{D}\hat{x}\exp\left(i\hat{x}T[x]-\frac{g^{2}}{2}\hat{x}C^{\phi}\hat{x}-\frac{1}{2}\phi\hat{C}^{\phi}\phi\right).$ (76)
F.1 Two-point functions
In the limit $N\to\infty$, the saddle point dominates this integral. The derivatives of the action, which are zero at the saddle point, are
@@ F.1 Two-point functions
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$. This recovers the same single-site process (Eq. 69, Eq. 70) and self-consistency condition (Eq. 71) described in Sec. E.
F.2 Four-point functions
@@
$Z=\exp\left(-N\mathcal{S}[C^{\phi}_{\star},0]\right)\int\mathcal{D}\delta C^{\phi}\mathcal{D}\delta\hat{C}^{\phi}\exp\left(-\frac{N}{2}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}^{T}\begin{pmatrix}H_{C^{\phi}C^{\phi}}&H_{C^{\phi}\hat{C}^{\phi}}\\ H_{\hat{C}^{\phi}{C}^{\phi}}^{T}&H_{\hat{C}^{\phi}\hat{C}^{\phi}}\end{pmatrix}\begin{pmatrix}\delta C^{\phi}\\ \delta\hat{C}^{\phi}\end{pmatrix}\right)$, (82)
@@
$N\left\langle\delta C^{\phi}(t_{1},t_{2})\delta C^{\phi}(t_{3},t_{4})\right\rangle_{\star}=\left[H_{C^{\phi}C^{\phi}}-H_{C^{\phi}\hat{C}^{\phi}}H_{\hat{C}^{\phi}\hat{C}^{\phi}}^{-1}H_{\hat{C}^{\phi}C^{\phi}}\right]^{-1}(t_{1},t_{2},t_{3},t_{4}).$
In principle, we have all the necessary components to evaluate this expression: the single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process. We could then take the temporal separation limits of Eq. 34 to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$. The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse. To circumvent this problem, we show in Appendix G that the temporal limits of Eq. 34 can be taken before taking the inverses. Under these temporal limits, the Hessian blocks are
@@
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$
@@
where we used the fact that multiplication by $i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$. Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space. In this representation, each Hessian block becomes a frequency-dependent scalar,
@@
$H_{C^{\phi}\hat{C}^{\phi}}(\bm{\omega})$ $=-1+g^{2}S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2}),$ (86)
@@
F.3 Adding sources